OpenStudy (anonymous):

solve the system and it would be a big help if you could tell me what the best system to use would be. x+3y=8 2x-y=2

OpenStudy (anonymous):

in this problem....I think substitution is the best x + 3y = 8 x = 8 - 3y now sub 8 - 3y in for x in the other equation 2x - y = 2 2(8 - 3y) - y = 2 16 - 6y - y = 2 16 - 7y = 2 -7y = 2 - 16 -7y = -14 y = -14/-7 y = 2 now sub 2 in for y in either of the original equations x + 3y = 8 x + 3(2) = 8 x + 6 = 8 x = 8 - 6 x = 2 check... 2x - y = 2 2(2) - 2 = 2 4 - 2 = 2 2 = 2 (correct) x = 2, y = 2 or (2,2) any questions ?

OpenStudy (anonymous):

wow no thanks a lot I think I got it now sorry I wans doing something \

OpenStudy (anonymous):

no problem...glad to help :)